OpenStudy (braves12):

help please g(x) = x + 1 and h(x) = x 2 - 4 h[g(x)] = _____.

OpenStudy (3mar):

May I help?

OpenStudy (braves12):

yes

OpenStudy (3mar):

Thanks

OpenStudy (3mar):

Do you know what does \[h[g(x)]\]mean?

OpenStudy (braves12):

no

OpenStudy (3mar):

That means : h is a function of the function g(x) i.e. replace every x in "h(x)" with the whole function "g(x)". Write it yourself and show me please.

OpenStudy (3mar):

@braves12

OpenStudy (mathmale):

Your " x 2 - 4" is vague. Did you mean "x^2 - 4?" Please clarify this. You are to find the "composition" of functions h(x) and g(x). Can you find examples of "composite functions" in your textbook or online reading materials?

OpenStudy (braves12):

yes it is "x^2 - 4 and no

OpenStudy (3mar):

So can you do what I mentioned above, braves12?

OpenStudy (braves12):

what do u mean

OpenStudy (3mar):

I mean replace every x in the function h(x) and put g(x), i.e (x+1)

OpenStudy (braves12):

g(x) = x 2 - 4

OpenStudy (3mar):

\[h(x)=x^2-4\] and \[g(x)\]

OpenStudy (3mar):

\[g(x)=x+1\]

OpenStudy (3mar):

\[h(x)=(g(x))^2-4=[(x+1)^2]-4\]

OpenStudy (3mar):

Got it?

OpenStudy (braves12):

yea kind of

OpenStudy (3mar):

Can you expand the parentheses and do math?

OpenStudy (braves12):

nope

OpenStudy (3mar):

\[h(x)=(x+1)^2-4=x^2+2x+1-4=x^2+2x\]

OpenStudy (braves12):

ok

OpenStudy (3mar):

sorry \[-3\]

OpenStudy (braves12):

what

OpenStudy (3mar):

I mean \[h(x)=x^2+2x-3\]

OpenStudy (3mar):

Did you get it?

OpenStudy (braves12):

nope

OpenStudy (3mar):

Where are you stuck?

OpenStudy (braves12):

when u expanded the parentheses

OpenStudy (3mar):

good!

OpenStudy (3mar):

\[h(x)=(x+1)^2-4\] start from here?

OpenStudy (braves12):

i dont know how to start from there

OpenStudy (3mar):

Welcome back!

OpenStudy (3mar):

@braves12

OpenStudy (braves12):

ok

OpenStudy (3mar):

so where you want to start?

OpenStudy (braves12):

idk

OpenStudy (3mar):

Do you have a Skype Id?

OpenStudy (braves12):

no

OpenStudy (3mar):

Ok

OpenStudy (3mar):

Let's start here

OpenStudy (3mar):

you know that g(x)=x+1 h(x)=x^2+4 ok?

OpenStudy (3mar):

Faster respond please!

OpenStudy (braves12):

yes

OpenStudy (3mar):

what if told you: find g(3) and h(4) for me Can you share your work?

OpenStudy (braves12):

i cant do that]

OpenStudy (3mar):

Where are you stuck?

OpenStudy (braves12):

all of it

OpenStudy (3mar):

For g(3): means replace every x in g(x) and put 3 instead!

OpenStudy (braves12):

WHAT

OpenStudy (3mar):

g(x) is a function of x we need to calculate this function at a certain point of its variable x so we want to determine g(x) at 3 Simple now?

OpenStudy (braves12):

kind of

OpenStudy (3mar):

Great! Happy for that! I need it 100%

OpenStudy (braves12):

yep

OpenStudy (3mar):

So, depend on your understanding can you believe that?\[g(3)=(3)+1=4\]

OpenStudy (braves12):

yes

OpenStudy (3mar):

Congratulations!

OpenStudy (3mar):

The same thing for h(x) ........ we need h(4) \[h(4)=(4)^2+4=16+4=20\] any problem???

OpenStudy (braves12):

nope

OpenStudy (3mar):

nope=don't understand? or nope=no problem?

OpenStudy (braves12):

no problem

OpenStudy (3mar):

Great! So we will do the same game for h(g(x)) instead we remove x in h(x) and put a NUMBER (3 OR 5 OR -1), we will remove it and put a FUNCTION [this function is g(x)]!

OpenStudy (braves12):

do what

OpenStudy (3mar):

are you asleep?

OpenStudy (braves12):

yea

OpenStudy (3mar):

yea about what? I was waiting and this is not a good evaluation for who are helping you!

OpenStudy (braves12):

that i was asleep

OpenStudy (3mar):

so another time!

OpenStudy (braves12):

ok

OpenStudy (3mar):

Salam!

OpenStudy (braves12):

what

OpenStudy (3mar):

Salam is our greetings, we are Muslims! Peace be upon you!